OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 4) with a slope of –1 in standard form.

jimthompson5910 (jim_thompson5910):

y - y1 = m(x - x1) y - y1= -1(x - x1) .. plug in the given slope y - 4= -1(x - (-1)) .. plug in the given point y - 4= -1(x + 1) now convert to standard form Ax+By = C. I'll let you do this

OpenStudy (anonymous):

1x=y=3?

jimthompson5910 (jim_thompson5910):

you wrote too many equal signs

OpenStudy (anonymous):

oppp I meant 1x+y=3

OpenStudy (anonymous):

look at the pic I attached for help

jimthompson5910 (jim_thompson5910):

you have the correct equation tabi_chick, nice job